OpenStudy (anonymous):

Please help! Alfonso is 8 years older than Ashley. Five years ago Alfonso was three times as old as Ashley. How old is Ashley now? So to solve this problem, I substituted Ashley for X and Alfonso for Y. I converted the word problem into 2 equations: y = x + 8 y -5 = 3x -5 Then I converted the first equation for x: x = 8 -y Then I solved for x...but something isn't right.

OpenStudy (strawberry17):

I set x = Alfanzo and y = Ashley x = y + 8 x - 5 = 3y Then I used substitution to solve the set:

OpenStudy (strawberry17):

I substituted the value of x from the first equation (y+8) into x in the second equation and simplified.

OpenStudy (strawberry17):

x - 5 = 3y (y + 8) - 5 = 3y y + 3 = 3y -y -y ---------- 3 = 2y 3 2y - = -- 2 2 3/2 = y

OpenStudy (anonymous):

oh ok! that makes sense. then y = 1.5

OpenStudy (strawberry17):

Yes! :)

OpenStudy (radar):

I know that the ages are 9 and 17 now, will try and show how.

OpenStudy (anonymous):

although...my textbook says that 1.5 is not the answer. It's higher: either 9, 17, or 8

OpenStudy (strawberry17):

hmm. okay, may radar can show how... to get right answer from textbook :)

OpenStudy (anonymous):

well thank you for coming over and making an effort =) @strawberry17

OpenStudy (strawberry17):

ur welcome! :)

OpenStudy (radar):

I am letting x = the age of the older one.Now then x - 8 = age of the youngest Now 5 years ago then x-5 is the older one 5 years ago. x-8-5 is the youngest one 5 years ago (x-13) we are told that the older one is now 3 times as old as the younger one (x-5) = 3(x-13) x-5 =3x -39 x-3x = -39 + 5 -2x= -34 x = 17 older one x-8=17 - 8 = 9 younger one

OpenStudy (radar):

Now using the convention you used: Alfonso is the older one so letting x = Alfonsos age now. The letting y = Ashley y= x - 8 or x = y + 8 Does that agree with your thinking?

OpenStudy (radar):

Now 5 years earlier: y -5 = 3(x-8-5) y-5 = 3x - 39. So far so good??

OpenStudy (anonymous):

yes but the younger one is x -5 so wouldn't that be: 17 -5?

OpenStudy (anonymous):

wait never mind...i looked back it's not like that

OpenStudy (radar):

No the older one is 17 and 17-5=12 is age of the olderr one 5yrs ago.

OpenStudy (radar):

the younger one would then be 9 - 5 or only 4 yrs old.

OpenStudy (anonymous):

ooooh ok. then the younger one is 9, the older one is 17

OpenStudy (anonymous):

thank you

OpenStudy (radar):

Yes, and I was going to attemp using your convention to solve it you using two unknowns x and y.

OpenStudy (radar):

But if you understand the way I worked it using only one variable x. I will leave.

OpenStudy (anonymous):

hhhhm? i don't understand...what you're wanting to do?

OpenStudy (anonymous):

you mean solve for y?

OpenStudy (radar):

I am at the point 5 years ago where y-5 = 3x -39

OpenStudy (anonymous):

ok. I wanted to find out "what age is Ashley now" so now that I've got that I'm done. but thanks for offering and thank you for helping me out =)

OpenStudy (radar):

O.K. You're method was legitimate, but somewhere you went awry and I was trying to see where.